# Maritime Cybersecurity: A Comprehensive Review

### v2

@@ 5.1.3. Blockchain-Based Defence
5.1.4. Network Level-Based Defence
@@
Maritime vessels are equipped with systems of varying criticality levels (Rajaram et al., 2022), with navigation and propulsion systems being mission-critical, in contrast to the non-essential nature of crew entertainment systems. A zone-based network segmentation approach, dividing ships’ network into isolated zones like the Global Ship Zone, Ship Control Zone, and Ship System Zone, is proposed in (Furumoto et al., 2020) that help contain potential threats and prevent them from spreading to other critical areas. Further, a cyber attack path discovery method (Polatidis et al., 2018) for maritime risk management can be used to identify potential vulnerabilities in network configurations.
@@
In (He et al., 2021), it was suggested that creating and maintaining whitelists of authorized devices and their permissions could provide a strong, regular security check for the local network environment of maritime IoT devices against DNS rebinding attacks.
5.1.5. Miscellaneous / Hybrid
@@
Security solutions for maritime threats do not fit neatly into the sole categories of machine/deep learning, blockchain, network level or cryptographic approaches but rather integrate various methods to achieve their goals. A rule-based risk model (Fossier, 2014) utilizing geospatial analysis, anomaly detection, and statistical methods is introduced to generate risk maps and real-time alerts, aimed at detecting abnormal behaviors indicative of maritime attacks. The ISOLA project (Laso et al., 2022) integrates a broad spectrum of technologies and methodologies, including sensor networks, data fusion, semantic reasoning, and visual analytics, while incorporating cyber vulnerability assessment tools and decision support systems to aid security personnel in identifying and responding to potential threats in the cruise ship industry. Chen and Wu (Chen and Wu, 2022) combined various technologies such as AIS, deep learning and cryptographic schemes to address multiple issues in maritime such as data integrity, authentication and network security.
@@
Table 4. The Evaluation of Existing Integrated/Holistic Countermeasures Against Most Common Threats Based on Maturity, Dataset, and Evaluation Method
Approach Asset Attacks Mitigated Maturity Dataset Evaluation Method
Pre-obtained Generated PoC Simulation Others
ML/DL (Kumar et al., 2021)
Man-in-the-Middle
Ransomware
DoS & DDoS
Backdoor Attacks
Injection Attacks
Experimental
(Ali et al., 2022)
DoS & DDoS
ARP Spoofing
Experimental
(Gyamfi et al., 2022)
DoS&DDoS
Network Intrusions
Experimental
(Wilson et al., 2024)
Attacks on Systems
Multi-stage Attacks
Experimental
Crypto (Chaudhry et al., 2021)
Unauthorised Access
GPS Spoofing
Message Tampering
Replay Attacks
Impersonation
Man-in-the-Middle
Theoretical
(Wang et al., 2021)
Replay Attacks
Data Breaches
Unauthorised Access
Identity spoofing
Man-in-the-Middle
Theoretical
(Liu et al., 2021)
Data Tampering
Unauthorised Access
Insider Threats
Data Unavailability
Man-in-the-Middle
Experimental
(Gupta et al., 2021)
Unauthorised Access
Data Leakage
Identity Spoofing
Man-in-the-Middle
Theoretical
(Hemminghaus et al., 2021)
Message Spoofing
Impersonation
Man-in-the-Middle
Experimental
Blockchain (Rahimi et al., 2020)
Impersonation
Data Integrity Attacks
Hijacking Attempts
GPS Spoofing
Experimental
(Zhang et al., 2022)
Replay Attacks
Information Leakage
Data Tampering
Experimental
(Gai et al., 2022)
Identity Disclosure
Privacy Breach
Unauthorised Access
Experimental
Network (Furumoto et al., 2020)
Malware Infections
CAN Protocol Attacks
Attacks on Systems
Theoretical
(Polatidis et al., 2018)
Lateral Movement
Privilege Escalation
Data Breaches
Experimental
(He et al., 2021)
DNS Rebinding
Device Impersonation
Data Exfiltration
Experimental
Hybrid (Fossier, 2014)
Vessel Abnormality
Interception Attempts
Experimental
(Laso et al., 2022)
Insider Threats
Physical Intrusions
Data Breaches
Theoretical
(Chen and Wu, 2022)
Radio Jamming
Privacy Breaches
DoS & DDoS
Experimental
Pre-obtained = Dataset already exists, and is acquired from an external source, rather than being created by the researcher themselves
@@ 5.2. Component Specific
Component-specific solutions refer to targeted countermeasures or protective measures designed to address vulnerabilities or prevent threats associated with particular vessel components. These solutions are tailored to the unique characteristics and vulnerabilities of each component, taking into account its function, exposure to threats, and critical role in the vessel’s overall operation. Component-specific solutions often involve specialized hardware, software, or operational procedures to provide deep, focused protection for critical components.
5.2.1. AIS Countermeasures
@@
Kowalska and Peel (Kowalska and Peel, 2012) presented an approach for detecting anomalous vessel behaviour using Gaussian Process (GP) models trained on AIS data. The GP anomaly detection method focuses on identifying suspicious vessel behaviors rather than securing the communication channel or data itself. Similarly, the work in (Ray et al., 2015) also focuses on identifying potential anomalies by processing incoming AIS data. Balduzzi et al. (Balduzzi et al., 2014) recommended applying anomaly detection techniques to the AIS data to identify suspicious activities, such as unexpected changes in a vessel’s route, which could indicate spoofing or hijacking attempts. The author (Balduzzi et al., 2014) also suggested to implement a Public Key Infrastructure (PKI) using X.509 certificates to enable authentication and integrity checks for AIS messages. Influenced by security solutions from IEEE 1609 standards for vehicle communications, Hall et al. (Hall et al., 2015) proposed a redesigned security protocol to address multiple security vulnerabilities in the current AIS vessel tracking system. To identify vessels that deliberately turn off their AIS transponders, Bernabé et al. (Bernabé et al., 2023) used a self-supervised deep learning approach to analyze patterns in AIS transmissions. By comparing these predictions to actual real-time observations, it can classify vessel trajectories as normal or potential suspicious. Iphar et al. (Iphar et al., 2015) outlined a conceptual methodology for using various data quality dimensions to assess the integrity and trustworthiness of information within AIS messages. Su et al. (Su et al., 2017) proposed authentication and privacy-preserving enhancements to the AIS protocol to address the security vulnerabilities that currently allow identity spoofing and enable tracking of vessels through their AIS broadcasts. They first introduced Digital Signature based Identity Authentication Scheme (DSIAS) to prevent tampering or forging of vessel identities and then proposed Mix-zone based Trajectory Privacy Protection Scheme (MTPPS) to provide anonymity of vessel trajectories. They further built a Blind-signature extension to the MTTPS, which provides an additional layer of privacy. Kontopoulos et al. (Kontopoulos et al., 2018) advocated an architectural solution for detecting malicious tampering of live AIS data streams. Goudossis and Katsikas (Goudossis and Katsikas, 2019) explored how a Maritime Certificate-less Identity-Based Public Key Cryptography (mIBC) infrastructure may enhance the security properties of AIS. While the previous paper introduced the concept of using Identity-Based cryptography to secure AIS, Goudossis and Katsikas (Goudossis and Katsikas, 2020) further delved into the implementation details, introduced a new application for seamless integration with the existing AIS infrastructure, and provided operational overhead estimates. Nguyen (Nguyen, 2020) evaluated the AIS architecture and explored using lightweight cryptographic algorithms to design an optimal authentication system. Sciancalepore et al. (Sciancalepore et al., 2021) focused on the design, security properties, implementation and performance evaluation of the proposed Auth-AIS protocol to secure vessel AIS broadcast messages in a practical manner. Kelly (Kelly, 2022) used Received Signal Strength Indicator (RSSI) based detection method to enhance AIS security by providing a means to identify and investigate vessels attempting to exploit the low power mode to evade tracking, which could indicate involvement in illicit activities. Deng et al. (Deng et al., 2023) proposed a novel lightweight Transformer-based network called GLFormer for specific emitter identification (SEI) to provide an extra security layer for AIS terminal emitters.
The maritime environment should necessitate caution when applying GPS research findings from other domains without targeted validation. Signal propagation patterns, interference sources, and operational requirements can differ substantially between land-based and maritime settings, potentially limiting the direct transferability of results. To assure the applicability and reliability of findings in GNSS threats and defenses, this study has prioritised the research that has conducted experiments and analyses in representative maritime conditions.
5.2.2. GNSS Countermeasures
@@
Grant et al. (Grant et al., 2009) tested the feasibility of eLoran as a backup system to provide PNT during GPS jamming incidents. Bhatti and Humphreys (Bhatti and Humphreys, 2017) proposed several countermeasures to mitigate the risk of GPS spoofing attacks on maritime vessels. In contrast, Drumhiller et al. (Drumhiller et al., 2017) focused on primary defenses against GPS jamming. Medina et al. (Medina et al., 2019) proposed jamming countermeasures like robust signal processing, adaptive antenna arrays, or multi-sensor fusion. A new technique for detecting GNSS spoofing attacks using signals from the Iridium satellite constellation was proposed by Oligeri et al. (Oligeri et al., 2020). Spravil et al. (Spravil et al., 2023) focused on detecting GPS spoofing attacks in the maritime domain using a novel software-based framework called MANA (MAritime Nmea-based Anomaly detection). Boudehenn et al. (Boudehenn et al., 2021) proposed a strategy to enhance the detection of navigation spoofing attacks and assess possible physical impacts.
5.2.3. ECDIS Countermeasures
@@
Based on the information provided by Svilicic et al. and Dyryavyy, some mitigation measures can be implemented to enhance the security of ECDIS systems. While the papers (Dyryavyy, 2014)(Svilicic et al., 2019a, b, c, 2020, d) offer valuable insights into ECDIS threats and propose mitigation measures, the authors did not include experiments specifically designed to test the effectiveness of those mitigation. These theoretical measures can be categorized into technical, procedural, and organizational aspects.
5.2.4. RADAR Countermeasures
@@
Various security solutions and mitigation measures have been developed to protect RADAR systems, both in general and for specific applications, although they are not exclusive to the maritime domain. These include deep learning-based anomaly detection techniques (Cohen et al., 2022)(de Riberolles et al., 2022), hash-based integrity checks and encryption designed for the ASTERIX protocol (Casanovas et al., 2015), and algorithmic approaches (Yang et al., 2018). Despite the existence of these security solutions, it can be challenging to directly apply them to the maritime scenarios (Wolsing et al., 2022). The unique characteristics and requirements of marine RADAR systems should necessitate tailored approaches to ensure effective security measures. Junior et al. (Leite Junior et al., 2021) briefly mentioned that as future work, they may develop tools to verify the integrity of the software used in naval RADAR systems, in order to detect any malicious code or malware that may have been pre-installed in the RADAR. Longo et al. (Longo et al., 2023c) proposed a detection system as a countermeasure to the marine RADAR attacks. It is designed as a policy enforcement system where the policies dictate how the RADAR should operate according to industry standards, regulations, and manufacturer specifications, in conjunction with onboard configurations. The experimental results demonstrate that the detection system proposed in this study consumes fewer resources while detecting these attacks with high accuracy.
5.2.5. VDR Countermeasures
@@
Söner et al. (Söner et al., 2023) also recommended several preventive and control measures to improve the cybersecurity of VDR based on their FMEA risk assessment findings. They emphasized that while cyberattacks cannot be entirely prevented due to the nature of the cyber world, their effects can be mitigated by conducting cyber risk assessments and implementing effective control measures to safeguard VDR from current and emerging cybersecurity threats. The authors (Hopcraft et al., 2023) also proposed several amendments to the standards to improve VDR security.
5.2.6. VSAT Countermeasures
@@
The author (Pavur et al., 2020) acknowledged that some of these solutions, such as end-to-end encryption, may have performance implications due to the high latency of satellite communications. However, they emphasized the need for the maritime industry to prioritize security and invest in the development of practical, satellite-optimized security measures to protect ships, crew, and cargo from potential cyber threats. Wu et al. (Wu et al., 2018) proposed a new lightweight authentication scheme called lite-CA (Lite Certification Authority) to reduce the amount of interaction required for authentication within the VSAT network architecture. To achieve real-time data encryption, they further proposed using a lightweight encryption algorithm called HW-F (high weight function) to replace traditional public key and symmetric encryption systems.
5.2.7. GMDSS Countermeasures
@@
Korcz (Korcz, 2023) discussed the ongoing modernization of GMDSS which includes introduction of new satellite providers beyond Inmarsat (e.g. Iridium and Beidou), development of the VHF Data Exchange System (VDES) for improved data communication and implementation of digital broadcasting of maritime safety information. These measures could provide more secure and efficient data exchange and enhance GMDSS communications. Osés and Juncadella (Martínez de Osés and Uyà Juncadella, 2021) argued that it is conceptually and technologically feasible to create a global VTS system as a counterpart to GMDSS.
@@
Table 5. The Evaluation of Existing Component Specific Countermeasures Against Most Common Threats Based on Maturity, Evaluation Method, and Performance
@@
AIS AIS Spoofing Drug Smuggling People Trafficking Trajectory Tracking Illegal Fishing AIS Message Injection AIS Message Flooding AIS Data Tampering Stream Poisoning Unusual Vessel Patterns Availability Disruption AIS Hijacking Detection model (Kowalska and Peel, 2012) Experimental
Detection technique (Ray et al., 2015) Theoretical $\mathord{?}?$
PKI using X.609 (Balduzzi et al., 2014) Theoretical $\mathord{?}?$
Redesigned protocol (Hall et al., 2015) Experimental
Transformer model (Bernabé et al., 2023) Experimental
Conceptual framework (Iphar et al., 2015) Theoretical $\mathord{?}?$
DSIAS and MTTPS (Su et al., 2017) Theoretical
Architectural solution (Kontopoulos et al., 2018) Experimental
mIBC (Goudossis and Katsikas, 2019)(Goudossis and Katsikas, 2020) Theoretical $\mathord{?}?$
Authenticator module (Nguyen, 2020) Experimental
Auth-AIS protocol (Sciancalepore et al., 2021) Experimental
Detection algorithm (Kelly, 2022) Experimental
Transformer model (Deng et al., 2023) Experimental
GNSS / GPS Intentional Jamming Interference Localised Jamming Wide-area Jamming GPS Service Denial GPS Spoofing Backup GPS (Grant et al., 2009) Experimental
Detection framework (Bhatti and Humphreys, 2017) Experimental
Improved equipment (Drumhiller et al., 2017) Theoretical $\mathord{?}?$
Multi-sensor fusion (Medina et al., 2019) Theoretical $\mathord{?}?$
Detection technique (Oligeri et al., 2020) Experimental
MANA framework (Spravil et al., 2023) Experimental
Detection system (Boudehenn et al., 2021) Experimental
ECDIS File Manipulation Directory Traversal Outdated Software Malware Introduction HTTP Header Injection Remote Code Execution Update system (Dyryavyy, 2014) Theoretical $\mathord{?}?$
Disable SMB v1 service (Svilicic et al., 2019a) Theoretical $\mathord{?}?$
Backup arrangement (Svilicic et al., 2019b) Theoretical $\mathord{?}?$
Proper configuration (Svilicic et al., 2019c) Theoretical $\mathord{?}?$
Access control (Svilicic et al., 2020) Theoretical $\mathord{?}?$
Secure setup (Svilicic et al., 2019d) Theoretical $\mathord{?}?$
RADAR Ship Trajectory Attack Network monitoring (Longo et al., 2023c) Experimental
VDR Ransomware Tampering Network segmentation (Söner et al., 2023) Theoretical $\mathord{?}?$
Physical measures (Hopcraft et al., 2023) Theoretical $\mathord{?}?$
VSAT Eavesdropping Spoofing Encryption method (Pavur et al., 2020) Theoretical $\mathord{?}?$
lite-CA (Wu et al., 2018) Experimental
GMDSS DoS Alert Disruption Improved quality (Korcz, 2023) Theoretical
Global VTS system (Martínez de Osés and Uyà Juncadella, 2021) Theoretical

### v3

@@ IV-A3 Blockchain-Based Defence
IV-A4 Network Level-Based Defence
@@
Maritime vessels are equipped with systems of varying criticality levels [124], with navigation and propulsion systems being mission-critical, in contrast to the non-essential nature of crew entertainment systems. A zone-based network segmentation approach, dividing ships’ network into isolated zones like the Global Ship Zone, Ship Control Zone, and Ship System Zone, is proposed in [125] that help contain potential threats and prevent them from spreading to other critical areas. Further, a cyber attack path discovery method [126] for maritime risk management can be used to identify potential vulnerabilities in network configurations.
@@
In [127], it was suggested that creating and maintaining whitelists of authorized devices and their permissions could provide a strong, regular security check for the local network environment of maritime IoT devices against DNS rebinding attacks.
IV-A5 Miscellaneous / Hybrid
@@
Security solutions for maritime threats do not fit neatly into the sole categories of machine/deep learning, blockchain, network level or cryptographic approaches but rather integrate various methods to achieve their goals. A rule-based risk model [128] utilizing geospatial analysis, anomaly detection, and statistical methods is introduced to generate risk maps and real-time alerts, aimed at detecting abnormal behaviors indicative of maritime attacks. The ISOLA project [129] integrates a broad spectrum of technologies and methodologies, including sensor networks, data fusion, semantic reasoning, and visual analytics, while incorporating cyber vulnerability assessment tools and decision support systems to aid security personnel in identifying and responding to potential threats in the cruise ship industry. Chen and Wu [130] combined various technologies such as AIS, deep learning and cryptographic schemes to address multiple issues in maritime such as data integrity, authentication and network security.
@@
TABLE III: The Evaluation of Existing Integrated/Holistic Countermeasures Against Most Common Threats Based on Maturity, Dataset, and Evaluation Method
Approach Asset Maturity Dataset Evaluation Method *Performance
Pre-obtained Generated Proof-of-Concept Simulation Others
ML/DL [112] Experimental
[113] Experimental
[114] Experimental
[115] Experimental
Crypto [116] Theoretical $\mathord{?}?$
[117] Theoretical $\mathord{?}?$
[118] Experimental
[119] Theoretical $\mathord{?}?$
[120] Experimental
Blockchain [121] Experimental
[122] Experimental
[123] Experimental
Network [125] Theoretical $\mathord{?}?$
[126] Experimental
[127] Experimental
Hybrid [128] Experimental
[129] Theoretical $\mathord{?}?$
[130] Experimental
Pre-obtained = Dataset already exists, and is not created by the researcher themselves
@@ IV-B Component Specific
Component-specific solutions refer to targeted countermeasures or protective measures designed to address vulnerabilities or prevent threats associated with particular vessel components. These solutions are tailored to the unique characteristics and vulnerabilities of each component, taking into account its function, exposure to threats, and critical role in the vessel’s overall operation. We have a summarized list here with additional comparative study of different approaches in Table IV.
IV-B1 AIS Countermeasures
@@
Kowalska and Peel [131] presented an approach for detecting anomalous vessel behaviour using Gaussian Process (GP) models trained on AIS data. The GP anomaly detection method focuses on identifying suspicious vessel behaviors rather than securing the communication channel or data itself. Similarly, the work in [132] also focuses on identifying potential anomalies by processing incoming AIS data. Balduzzi et al. [85] recommended applying anomaly detection techniques to the AIS data to identify suspicious activities, such as unexpected changes in a vessel’s route, which could indicate spoofing or hijacking attempts. The author [85] also suggested to implement a Public Key Infrastructure (PKI) using X.509 certificates to enable authentication and integrity checks for AIS messages. Influenced by security solutions from IEEE 1609 standards for vehicle communications, Hall et al. [133] proposed a redesigned security protocol to address multiple security vulnerabilities in the current AIS vessel tracking system. To identify vessels that deliberately turn off their AIS transponders, Bernabé et al. [134] used a self-supervised deep learning approach to analyze patterns in AIS transmissions. By comparing these predictions to actual real-time observations, it can classify vessel trajectories as normal or potential suspicious. Iphar et al. [135] outlined a conceptual methodology for using various data quality dimensions to assess the integrity and trustworthiness of information within AIS messages. Su et al. [136] proposed authentication and privacy-preserving enhancements to the AIS protocol to address the security vulnerabilities that currently allow identity spoofing and enable tracking of vessels through their AIS broadcasts. They first introduced Digital Signature based Identity Authentication Scheme (DSIAS) to prevent tampering or forging of vessel identities and then proposed Mix-zone based Trajectory Privacy Protection Scheme (MTPPS) to provide anonymity of vessel trajectories. They further built a Blind-signature extension to the MTTPS, which provides an additional layer of privacy. Kontopoulos et al. [137] advocated an architectural solution for detecting malicious tampering of live AIS data streams. Goudossis and Katsikas [138] explored how a Maritime Certificate-less Identity-Based Public Key Cryptography (mIBC) infrastructure may enhance the security properties of AIS. While the previous paper introduced the concept of using Identity-Based cryptography to secure AIS, Goudossis and Katsikas [139] further delved into the implementation details, introduced a new application for seamless integration with the existing AIS infrastructure, and provided operational overhead estimates. Nguyen [140] evaluated the AIS architecture and explored using lightweight cryptographic algorithms to design an optimal authentication system. Sciancalepore et al. [141] focused on the design, security properties, implementation and performance evaluation of the proposed Auth-AIS protocol to secure vessel AIS broadcast messages in a practical manner. Kelly [142] used Received Signal Strength Indicator (RSSI) based detection method to enhance AIS security by providing a means to identify and investigate vessels attempting to exploit the low power mode to evade tracking, which could indicate involvement in illicit activities. Deng et al. [143] proposed a novel lightweight Transformer-based network called GLFormer for specific emitter identification (SEI) to provide an extra security layer for AIS terminal emitters.
IV-B2 GNSS Countermeasures
@@
Grant et al. [100] tested the feasibility of eLoran as a backup system to provide PNT during GPS jamming incidents. Bhatti and Humphreys [101] proposed several countermeasures to mitigate the risk of GPS spoofing attacks on maritime vessels. In contrast, Drumhiller et al. [144] focused on primary defenses against GPS jamming. Medina et al. [102] proposed jamming countermeasures like robust signal processing, adaptive antenna arrays, or multi-sensor fusion. A new technique for detecting GNSS spoofing attacks using signals from the Iridium satellite constellation was proposed by Oligeri et al. [145]. Spravil et al. [146] focused on detecting GPS spoofing attacks in the maritime domain using a novel software-based framework called MANA (MAritime Nmea-based Anomaly detection). Boudehenn et al. [147] proposed a strategy to enhance the detection of navigation spoofing attacks and assess possible physical impacts.
IV-B3 ECDIS Countermeasures
@@
Based on the information provided by Svilicic et al. and Dyryavyy, some mitigation measures can be implemented to enhance the security of ECDIS systems. While the papers [108][103, 104, 105, 106, 107] offer valuable insights into ECDIS threats and propose mitigation measures, the authors did not include experiments specifically designed to test the effectiveness of those mitigation. These theoretical measures can be categorized into technical, procedural, and organizational aspects.
IV-B4 RADAR Countermeasures
@@
Various security solutions and mitigation measures have been developed to protect RADAR systems, both in general and for specific applications, although they are not exclusive to the maritime domain. These include deep learning-based anomaly detection techniques [148][149], hash-based integrity checks and encryption designed for the ASTERIX protocol [150], and algorithmic approaches [151]. Despite the existence of these security solutions, it can be challenging to directly apply them to the maritime scenarios [95]. The unique characteristics and requirements of marine RADAR systems should necessitate tailored approaches to ensure effective security measures. Junior et al. [94] briefly mentioned that as future work, they may develop tools to verify the integrity of the software used in naval RADAR systems, in order to detect any malicious code or malware that may have been pre-installed in the RADAR. Longo et al. [96] proposed a detection system as a countermeasure to the marine RADAR attacks. It is designed as a policy enforcement system where the policies dictate how the RADAR should operate according to industry standards, regulations, and manufacturer specifications, in conjunction with onboard configurations. The experimental results demonstrate that the detection system proposed in this study consumes fewer resources while detecting these attacks with high accuracy.
IV-B5 VDR Countermeasures
@@
Söner et al. [93] also recommended several preventive and control measures to improve the cybersecurity of VDR based on their FMEA risk assessment findings. They emphasized that while cyberattacks cannot be entirely prevented due to the nature of the cyber world, their effects can be mitigated by conducting cyber risk assessments and implementing effective control measures to safeguard VDR from current and emerging cybersecurity threats. The authors [92] also proposed several amendments to the standards to improve VDR security.
IV-B6 VSAT Countermeasures
@@
The author [99] acknowledged that some of these solutions, such as end-to-end encryption, may have performance implications due to the high latency of satellite communications. However, they emphasized the need for the maritime industry to prioritize security and invest in the development of practical, satellite-optimized security measures to protect ships, crew, and cargo from potential cyber threats. Wu et al. [152] proposed a new lightweight authentication scheme called lite-CA (Lite Certification Authority) to reduce the amount of interaction required for authentication within the VSAT network architecture. To achieve real-time data encryption, they further proposed using a lightweight encryption algorithm called HW-F (high weight function) to replace traditional public key and symmetric encryption systems.
IV-B7 GMDSS Countermeasures
@@
Korcz [153] discussed the ongoing modernization of GMDSS which includes introduction of new satellite providers beyond Inmarsat (e.g. Iridium and Beidou), development of the VHF Data Exchange System (VDES) for improved data communication and implementation of digital broadcasting of maritime safety information. These measures could provide more secure and efficient data exchange and enhance GMDSS communications. Osés and Juncadella [154] argued that it is conceptually and technologically feasible to create a global VTS system as a counterpart to GMDSS.
@@
TABLE IV: The Evaluation of Existing Component Specific Countermeasures Against Most Common Threats Based on Maturity, Evaluation Method, and Performance
@@
AIS AIS Spoofing Drug Smuggling People Trafficking Trajectory Tracking Illegal Fishing AIS Message Injection AIS Message Flooding AIS Data Tampering Stream Poisoning Unusual Vessel Patterns Availability Disruption AIS Hijacking Detection model [131] Experimental
Detection technique [132] Theoretical $\mathord{?}?$
PKI using X.609 [85] Theoretical $\mathord{?}?$
Redesigned protocol [133] Experimental
Transformer model [134] Experimental
Conceptual framework [135] Theoretical $\mathord{?}?$
DSIAS and MTTPS [136] Theoretical
Architectural solution [137] Experimental
mIBC [138][139] Theoretical $\mathord{?}?$
Authenticator module [140] Experimental
Auth-AIS protocol [141] Experimental
Detection algorithm [142] Experimental
Transformer model [143] Experimental
GNSS / GPS Intentional Jamming Interference Localised Jamming Wide-area Jamming GPS Service Denial GPS Spoofing Backup GPS [100] Experimental
Detection framework [101] Experimental
Improved equipment [144] Theoretical $\mathord{?}?$
Multi-sensor fusion [102] Theoretical $\mathord{?}?$
Detection technique [145] Experimental
MANA framework [146] Experimental
Detection system [147] Experimental
ECDIS File Manipulation Directory Traversal Outdated Software Malware Introduction HTTP Header Injection Remote Code Execution Update system [108] Theoretical $\mathord{?}?$
Disable SMB v1 service [103] Theoretical $\mathord{?}?$
Backup arrangement [104] Theoretical $\mathord{?}?$
Proper configuration [105] Theoretical $\mathord{?}?$
Access control [106] Theoretical $\mathord{?}?$
Secure setup [107] Theoretical $\mathord{?}?$
RADAR Ship Trajectory Attack Network monitoring [96] Experimental
VDR Ransomware Tampering Network segmentation [93] Theoretical $\mathord{?}?$
Physical measures [92] Theoretical $\mathord{?}?$
VSAT Eavesdropping Spoofing Encryption method [99] Theoretical $\mathord{?}?$
lite-CA [152] Experimental
GMDSS DoS Alert Disruption Improved quality [153] Theoretical
Global VTS system [154] Theoretical
V Challenge Discussions